OpenStudy (anonymous):

Can anyone help me? sin(x + pi/2) = cosx How do I prove this??

OpenStudy (solomonzelman):

First thing that pops up n my head. \(\large\color{blue}{ \bf sin(x+90)=cos(x) }\) \(\large\color{blue}{ \bf sin(x)cos(90)+sin(90)cos(x)=cos(x) }\) \(\large\color{blue}{ \bf sin(x)~(0)+(1)~cos(x)=cos(x) }\) ....

OpenStudy (solomonzelman):

Just using the sin(a+b) identity

OpenStudy (anonymous):

Use the sum formula for sine to simplify the expression. The formula states that sin(A+B)=sinAcosB+cosAsinB. sin(x)cos(π2)+cos(x)sin(π2) Take the cosine of π2 to get 0. sin(x)(0)+cos(x)sin(π2) Remove the parentheses. sin(x)⋅0+cos(x)sin(π2) Multiply sin(x) by 0 to get 0. (0)+cos(x)sin(π2) Remove the parentheses. 0+cos(x)sin(π2) Take the sine of π2 to get 1. 0+cos(x)(1) Remove the parentheses. 0+cos(x)⋅1 Multiply cos(x) by 1 to get cos(x). 0(cos(x)) Remove the parentheses. 0+cos(x) Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. cos(x)

OpenStudy (anonymous):

Oh okay thanks! I guess I just messed up my formula...